OpenStudy (anonymous):

please help me : how to solve g(x)= x^2 +6x +1

OpenStudy (luigi0210):

You could use the quadratic formula?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

I can change it into standard form, I just don't know how to solve it

OpenStudy (luigi0210):

Well you could solve it by 3 methods, factoring, quadratic formula, or graphing. That equation doesn't seem factorable to me

OpenStudy (anonymous):

But how do I solve it in the quadratic formula?

OpenStudy (luigi0210):

Use: \(\Large \frac{b \pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (luigi0210):

From \(\Large f(x)= Ax^2+Bx+C\)

OpenStudy (anonymous):

thank you so much!!